# bibliography.bib

### v1

@@ Abstract
We analyze the near-horizon symmetries of static, axisymmetric, four-dimensional black holes with spherical and toroidal horizon topologies in vacuum general relativity. These black hole solutions, collectively referred to as local/distorted black holes, are known in closed form and are not asymptotically flat. Building on earlier works in the literature that primarily focused on black holes with spherical topology, we compute the algebra of the Killing vector fields that preserve the asymptotic structure near the horizons and the algebra of the associated Noether-Wald charges under the boundary conditions that produce the spin-$s$ BMSd and the Heisenberg-like algebras. We show that a similar analysis extends to all local axisymmetric black holes. The toroidal topology of the holes changes the algebras considerably. For example, one obtains two copies of spin-$s$ BMS3 instead of spin-$s$ BMS4. We also revisit the thermodynamics of black holes under these boundary conditions. While previous studies suggested that spin-$s$ BMSd preserves the first law of thermodynamics for isolated horizons ($\kappa=\text{const.}$.), our analysis indicates that this is not generally the case when the spin parameter $s$ is nonzero. A nonzero $s$ can be seen as introducing a conical singularity (in the Euclidean quantum gravity sense) or a Hamiltonian that causes soft hairs to contribute to the energy. This leads us to interpret the spin-$s$ BMSd boundary condition as arising in the context of dynamical black holes.
@@ 1 Introduction
In 1962, Bondi, van der Burg, Metzner, and Sachs, while analyzing gravitational waves in asymptotically flat spacetimes, discovered that the coordinate transformations that preserve the asymptotic structure of the metric near future null infinity $(\mathscr{I}^{+})$ formed a symmetry group much larger than the Poincaré group, now known as the Bondi-Metzner-Sachs (BMS) group [bondi1962, sachs1962_waves, sachs1962_asym]. It is a semi-direct product of the Lorentz group and an infinite-dimensional group of transformations called “supertranslations”. Around the same time, Penrose, Newman, Geroch, and others developed a topological approach to asymptotically flat spacetimes [Penrose:1962, Penrose:1965, Newman-Penrose-1968, Geroch1977]. These efforts allowed the BMS group to be reinterpreted as the exact global diffeomorphism group of $\mathscr{I}^{+}$ (after quotienting out the trivial diffeomorphisms, see, for example, [ashtekar2014]). On the quantum side, since the BMS group only admitted discrete-spin (irreducible) representations, in contrast to the Poincaré group which also allows continuous-spin representations, it offered an explanation for why nature favors discrete-spin particles, including the spin-$\frac{1}{2}$ particles [Newman1965, mccarthy1972_structure, McCarthy-74].
@@
After decades of relative obscurity, the BMS group has experienced a resurgence of interest in recent years, driven by the equivalences between gravitational memory effect, soft theorems, and asymptotic symmetries [Strominger-Zhiboedov:2014]. It has also been found that the Lorentz part of the BMS group can be expanded to a larger group of “superrotations” [Barnich_2011]. The associated Noether charges form a Witt algebra, a special case of the Virasoro algebra with no central extension, while the supertranslations produce an Abelian current [Barnich_2011]. Generally, by “BMS group” these days, we mean the group including superrotations.
[…]
@@
In this paper, we follow the above approaches to study the near-horizon symmetries of static, axisymmetric (Weyl) black holes in vacuum general relativity. These black holes, distorted by matter sources at a distance, are not asymptotically flat (except for the special case of the Schwarzschild solution) and can have spherical or toroidal horizon topologies despite being vacuum solutions [stephani_exact_sol]. In their seminal work, Geroch and Hartle obtained these solutions in closed form and referred to them as “local black holes” [GerochHartle]. Chandrasekhar later re-derived the spherical topology solutions in Schwarzschild-like coordinates, demonstrating how these can be seen as distortions of the Schwarzschild metric by distant multipole sources [Doroshkevich-Zeldovich-65, chandrasekharbook]. Xanthopoulos extended Chandrasekhar’s treatment to toroidal holes [LocalToroidalXanthapolous].
@@ 2 General Treatment
In studies of asymptotic symmetries in general relativity, null coordinates are often employed. While earlier works on near-horizon symmetries utilized null coordinates [Donnay:2015, Donnay:2016], finding them in closed form is challenging for non-asymptotically flat spacetimes or horizons with non-spherical topology, such as those we considered in this paper. We thus adopt Gaussian-type coordinates (that is, $\delta_{\mu\rho}$ to first order) as used in later works [Donnay:2019, Giribet2023, Grumiller-2020] — these works take inspiration from the “membrane paradigm” of black hole horizons pioneered by Price and Thorne [PriceThorne]. We would like to add here that earlier works also considered near-horizon symmetries in the context of horizon microstates [Carlip_1995, Carlip_1999, Hotta_2001, Carlip:2002].
In this section, we summarize the results and approaches adopted in [Donnay:2016, Grumiller-2020], presenting the necessary details in a self-contained manner for use in the subsequent sections. In either approach, one begins by finding a suitable radial coordinate and writes the black-hole geometry as a product of horizon geometry and a codimension-two Minkowski metric in Rindler coordinates. This is similar to the approach one takes in Euclidean quantum gravity where, in addition, one periodically identifies the Euclideanized time, and the periodicity needed to obtain a non-conical two-dimensional Euclidean space gives the temperature of the hole (see, for example, [Gibbons-Hawking-Book]). If $\rho$ is the suitable radial coordinate such that the horizon is located at $\rho=0$, one can impose the following boundary conditions [Grumiller-2020] [^4] These fall-off conditions are a bit more general than those used in [Donnay:2016]. However, the Lie algebra of the Killing vectors will not differ between the two fall-offs.
@@
where $T$ and $\Phi^{A}$ are related via $\partial_{t}T+\Phi^{A}\partial_{A}\kappa=\delta\kappa$ (which follows from the fall-off conditions on Lie derivatives of the metric, see equation (2.12) in [Donnay:2016]), but are otherwise arbitrary functions of their arguments. For constant $\kappa$, this equation implies that $T$ is only a function of the angular coordinates. In any case, $T$ and $\Phi^{A}$ represent supertranslations and superrotations, respectively.
@@
Killing-Vector Algebra: To find the algebra of the Killing vectors, one makes use of the modified Lie bracket [Barnich-Troessaert]
$[\xi_{1},\xi_{2}]=\mathcal{L}_{\xi_{1}}\xi_{2}-\delta_{\xi_{1}}\xi_{2}+\delta_{\xi_{2}}\xi_{1}~{},$ (3)
where, $\delta_{\xi_{i}}\xi_{j}$ represents the change induced in $\xi_{j}[g]$ by the variation $\mathcal{L}_{\xi_{i}}g_{\mu\nu}$, that is, $\delta_{\xi_{1}}\xi_{2}^{\mu}=\xi_{2}^{\mu}[\mathcal{L}_{\xi_{1}}g]$ The (closed) algebra then works out to be
$[\xi(T_{1},\Phi_{1}^{A}),\xi(T_{2},\Phi_{2}^{A})]^{\mu}=\xi_{1}^{\nu}\partial_{\nu}\xi_{2}^{\mu}-\xi_{2}^{\nu}\partial_{\nu}\xi_{1}^{\mu}-\xi_{2}^{\mu}[\mathcal{L}_{\xi_{1}}g]+\xi_{1}^{\mu}[\mathcal{L}_{\xi_{2}}g]~{},$ (4)
@@
For spherical black holes and constant surface gravity $\kappa$ the algebra has been worked out [Donnay:2016, Donnay:2019]. One uses the complex coordinates $\zeta,\bar{\zeta}$ on the sphere, expands $T(x^{A})$ and $\Phi^{A}(x^{A})$ in Laurent modes
@@
This is a semi-direct sum of two commuting copies of the Witt algebra generated by $\mathcal{Y}_{m}$ and $\bar{\mathcal{Y}}_{n}$ which represent the superrotations, and an Abelian current algebra generated by $\mathcal{T}_{mn}$’s which represent the supertranslations. This, for example, has been verified for various C-metrics at both their black hole and acceleration horizons as well as for the (cosmological) de Sitter horizon [Donnay:2019, Anabalon2021]. One can physically interpret the mode $\mathcal{T}_{00}$ as the generator of rigid translations and therefore can be associated with energy. All other generators, i.e., $\mathcal{Y}_{m}$, $\bar{\mathcal{Y}}_{n}$ and $\mathcal{T}_{mn}$, commute with $\mathcal{T}_{00}$, and they can thus be generators of soft hair on the horizon [Donnay:2016].
Charge Algebras: For the charge algebra, one starts with the charge variation, which can be obtained from the covariant formalism [Regge_Teitelboim_1974, Wald-Zoupas, Barnich_Brandt_2002]
[…]
where $\left(\mathrm{d}^{d-2}x\right)_{\mu\nu}\equiv\frac{1}{(d-2)!}\varepsilon_{\mu\nu\alpha_{1}\alpha_{2}\cdots\alpha_{d-2}}\mathrm{d}x^{\alpha_{1}}\wedge\mathrm{d}x^{\alpha_{2}}\wedge\cdots\wedge\mathrm{d}x^{\alpha_{d-2}}$, $h_{\mu\nu}\equiv\delta g_{\mu\nu}$, $\xi$ is the asymptotic Killing vector, and the slash on $\delta$ represents that the variation might not be integrable. Using the boundary conditions (1), and the Killing vectors (2), one obtains the following dynamical fields [Grumiller-2020]:
@@
The integrability of $\delta Q$ will depend on the boundary conditions imposed. This is integrable, for example, for constant surface gravity [Donnay:2016, Donnay:2019]. The Poisson brackets of the charges in this case work out to be a semi-direct sum of two copies of the Witt and an Abelian current algebra, just the same as their Killing vector algebra above in (LABEL:WittAbelian) [Donnay:2015, Donnay:2016, Donnay:2019]. This is very close to the BMS (with superrotations) algebra with slightly different structure constants. Note that the expression for variation of charges in [Donnay:2016, Donnay:2019] (see, for example, equation (2.32) in [Donnay:2016]) looks slightly different from equation (12). However, they are identical expressions in different notations.
Considering the problem in a rotating frame within the Arnowitt–Deser–Misner (ADM) framework of general relativity, one can introduce a “spin parameter” $s$ and impose suitable boundary conditions to ensure that the boundary term — typically required for a well-defined Hamiltonian formulation — is integrable. Under these conditions, a deformation of the BMS group, known as the “spin-$s$ BMS group,” emerges [Grumiller-2020] (see Figure 1). For $s=0$, this recovers the algebra obtained by [Donnay:2016, Donnay:2019], described above in (LABEL:WittAbelian), while for $s=1$, it yields the BMSd algebra.
@@
A second set of boundary conditions (without the spin parameter) leads to the Heisenberg algebra for the charge generators [Grumiller-2020]. The boundary conditions for obtaining the BMS and the Heisenberg algebra are detailed below. For more details, see [Grumiller-2020, Grumiller-Jabbari-Book]. Boundary Condition for BMS: The ADM formalism is characterized by the lapse function $N(x^{\mu})$ and the shift vector $N^{i}(x^{\mu})$. Near a horizon they can be expanded as follows [Grumiller-2020]
@@
This factor is crucial in obtaining the spin-$s$ BMS algebra. We will return to this point in section 4 when discussing thermodynamics. The total Hamiltonian is the generator of unit time translations, given by [Grumiller-2020]
@@
The Poisson brackets of the charges work out to be [Grumiller-2020]
@@
This algebra is the semi-direct sum of diffeomorphisms at the spacelike section of the horizon generated by $\mathcal{J}_{A}$, and an Abelian algebra generated by $\mathcal{P}^{(s)}$. As before, introducing complex coordinates on the sphere, [Grumiller-2020] expanded supertranslations and superrotations in Laurent modes. The corresponding generators can be defined as
@@
whose algebra works out to be [Grumiller-2020]
$\begin{gathered}\{\mathcal{J}_{k},\mathcal{P}^{(s)}_{(m,n)}\}=\left(\frac{s}{2}(k+1)-m\right)\mathcal{P}^{(s)}_{(k+m,n)}~{},\qquad\{\bar{\mathcal{J}}_{k},\mathcal{P}^{(s)}_{(m,n)}\}=\left(\frac{s}{2}(k+1)-n\right)\,\mathcal{P}^{(s)}_{(m,k+n)}\\ \{\mathcal{J}_{n},\mathcal{J}_{m}\}=(n-m)\,\mathcal{J}_{n+m}~{},\quad\{\bar{\mathcal{J}}_{n},\bar{\mathcal{J}}_{m}\}=(n-m)\,\bar{\mathcal{J}}_{n+m}~{},\\ \{\mathcal{J}_{n},\,\bar{\mathcal{J}}_{m}\}=\{\mathcal{P}^{(s)}_{(m,n)},\,\mathcal{P}^{(s)}_{(m^{\prime},n^{\prime})}\}=0~{}.\end{gathered}$ (23)
The third and the fourth brackets — $\{\mathcal{J}_{n},\mathcal{J}_{m}\}$ and $\{\bar{\mathcal{J}}_{n},\bar{\mathcal{J}}_{m}\}$ — denote two copies of the Witt algebra. This is the spin-$s$ BMSd algebra. For $s=0$, this is the algebra for constant surface gravity in [Donnay:2015, Donnay:2016], which is the Poisson bracket counterpart of (LABEL:WittAbelian). The different BMS groups and their relations are shown in Figure 1.
Figure 1: Evolution of the BMS group: Originally discovered by Bondi, Metzner, van der Burg, and Sachs, the BMS group was later generalized in 2011 by Barnich and Troessaert to include superrotations. Various deformations of the BMS algebra were subsequently studied (see [Safari-Jabbari:2019] and references therein).
@@
One gets spin-$s$ BMS3 algebra [Barnich:2006, Grumiller-2020]:
@@
Boundary Conditions for Heisenberg-like Algebra: Algebra depends on the choice of slicing used in the solution space [Grumiller-Jabbari-Book]. Interchanging the roles of charges $(\mathcal{P},\mathcal{J}_{A})$ and vectors $(T,\Phi^{A})$ in the solution space
@@
the charge variation can also be integrated [Grumiller-2020]. Define $\mathcal{J}^{H}_{A}=\mathcal{J}_{A}\mathcal{P}^{-1}$. Then the charges (13) take the form
@@
When $F_{AB}=0$, this is exactly Heisenberg algebra [Grumiller-2020].

### v2

# Near-Horizon Symmetries of Local Black Holes in General Relativity

@@ Abstract
We analyze the near-horizon symmetries of static, axisymmetric, four-dimensional black holes with spherical and toroidal horizon topologies in vacuum general relativity. These black hole solutions, collectively referred to as local/distorted black holes, are known in closed form and are not asymptotically flat. Building on earlier works in the literature that primarily focused on black holes with spherical topology, we compute the algebra of the Killing vector fields that preserve the asymptotic structure near the horizons and the algebra of the associated Noether-Wald charges under the boundary conditions that produce the spin-$s$ BMSd and the Heisenberg-like algebras. We show that a similar analysis extends to all local axisymmetric black holes. The toroidal topology of the holes changes the algebras considerably. For example, one obtains two copies of spin-$s$ BMS3 instead of spin-$s$ BMS4. We also revisit the thermodynamics of black holes under these boundary conditions. While previous studies suggested that spin-$s$ BMSd preserves the first law of thermodynamics for isolated horizons ($\kappa=\text{const.}$), our analysis indicates that this is not generally the case when the spin parameter $s$ is nonzero. A nonzero $s$ can be seen as introducing a conical singularity (in the Euclidean quantum gravity sense) or a Hamiltonian that causes soft hairs to contribute to the energy. This leads us to interpret the spin-$s$ BMSd boundary condition as arising in the context of dynamical black holes.
@@ 1 Introduction
In 1962, Bondi, van der Burg, Metzner, and Sachs, while analyzing gravitational waves in asymptotically flat spacetimes, discovered that the coordinate transformations that preserve the asymptotic structure of the metric near future null infinity $(\mathscr{I}^{+})$ formed a symmetry group much larger than the Poincaré group, now known as the Bondi-Metzner-Sachs (BMS) group [1, 2, 3]. It is a semi-direct product of the Lorentz group and an infinite-dimensional group of transformations called “supertranslations”. Around the same time, Penrose, Newman, Geroch, and others developed a topological approach to asymptotically flat spacetimes [4, 5, 6, 7]. These efforts allowed the BMS group to be reinterpreted as the exact global diffeomorphism group of $\mathscr{I}^{+}$ (after quotienting out the trivial diffeomorphisms, see, for example, [8]). On the quantum side, since the BMS group only admitted discrete-spin (irreducible) representations, in contrast to the Poincaré group which also allows continuous-spin representations, it offered an explanation for why nature favors discrete-spin particles, including the spin-$\frac{1}{2}$ particles [9, 10, 11].
@@
After decades of relative obscurity, the BMS group has experienced a resurgence of interest in recent years, driven by the equivalences between gravitational memory effect, soft theorems, and asymptotic symmetries [12]. It has also been found that the Lorentz part of the BMS group can be expanded to a larger group of “superrotations” [13]. The associated Noether charges form a Witt algebra, a special case of the Virasoro algebra with no central extension, while the supertranslations produce an Abelian current [13]. Generally, by “BMS group” these days, we mean the group including superrotations.
In 2015, Hawking realized that supertranslations near the horizons of stationary black holes could provide a possible resolution to the information loss paradox [14] and the zero-energy conserved charges or “soft hairs” associated with these symmetries could encode information on the black hole horizons, as a form of holography [15, 16]. Subsequently, a number of authors have studied the near-horizon symmetries of various black holes and worked out the associated algebras [17, 18, 19, 20, 21, 22]. In particular, with suitable fall-off conditions, the symmetry algebra (for spherical black holes) was found to be a semi-direct sum of two copies of the Witt algebra and an Abelian current algebra with structure constants slightly different from those of the BMS algebra [17, 18]. Considering the fall-off conditions in a $3+1$ Arnowitt-Deser-Misner (ADM) decomposition, it was subsequently shown that one can obtain “spin-$s$ BMSd algebra” for a certain choice of boundary condition parametrized by a parameter $s$ [20]. This algebra gives the familiar BMSd algebra for s=1s=1444In Figure 1, we show the different BMS groups and the relations among themselves and the Poincaré group.. For another boundary condition, one obtains a Heisenberg-like algebra, the generators of which reproduce the generators of the BMS-like algebra under composition [23, 20]. Similar infinite-dimensional algebras are also found near generic null hypersurfaces (see, for example, [24, 25, 26]).
@@
In this paper, we follow the above approaches to study the near-horizon symmetries of static, axisymmetric (Weyl) black holes in vacuum general relativity. These black holes, distorted by matter sources at a distance, are not asymptotically flat (except for the special case of the Schwarzschild solution) and can have spherical or toroidal horizon topologies despite being vacuum solutions [27]. In their seminal work, Geroch and Hartle obtained these solutions in closed form and referred to them as “local black holes” [28]. Chandrasekhar later re-derived the spherical topology solutions in Schwarzschild-like coordinates, demonstrating how these can be seen as distortions of the Schwarzschild metric by distant multipole sources [29, 30]. Xanthopoulos extended Chandrasekhar’s treatment to toroidal holes [31].
@@ 2 General Treatment
In studies of asymptotic symmetries in general relativity, null coordinates are often employed. While earlier works on near-horizon symmetries utilized null coordinates [17, 18], finding them in closed form is challenging for non-asymptotically flat spacetimes or horizons with non-spherical topology, such as those we consider in this paper. We thus adopt Gaussian-type coordinates (that is, $\delta_{\mu\rho}$ to first order) as used in later works [19, 22, 20] — these works take inspiration from the “membrane paradigm” of black hole horizons pioneered by Price and Thorne [32]. We would like to add here that some earlier works also considered near-horizon symmetries in the context of horizon microstates [33, 34, 35, 36].
In this section, we summarize the results and approaches adopted in [18, 20], presenting the necessary details in a self-contained manner for use in the subsequent sections. In either approach, one begins by finding a suitable radial coordinate and writes the black-hole geometry as a product of horizon geometry and a codimension-two Minkowski metric in Rindler coordinates. This is similar to the approach one takes in Euclidean quantum gravity where, in addition, one periodically identifies the Euclideanized time, and the periodicity needed to obtain a non-conical two-dimensional Euclidean space gives the temperature of the hole (see, for example, [37]). If $\rho$ is the suitable radial coordinate such that the horizon is located at $\rho=0$, one can impose the following boundary conditions [20] [^5] These fall-off conditions are a bit more general than those used in [18]. However, the Lie algebra of the Killing vectors will not differ between the two fall-offs.
@@
where $T$ and $\Phi^{A}$ are related via $\partial_{t}T+\Phi^{A}\partial_{A}\kappa=\delta\kappa$ (which follows from the fall-off conditions on Lie derivatives of the metric, see equation (2.12) in [18]), but are otherwise arbitrary functions of their arguments. For constant $\kappa$, this equation implies that $T$ is only a function of the angular coordinates. In any case, $T$ and $\Phi^{A}$ represent supertranslations and superrotations, respectively.
@@
Killing-Vector Algebra: To find the algebra of the Killing vectors, one makes use of the modified Lie bracket [38]
$\xi_{1},\xi_{2}]=\mathcal{L}_{\xi_{1}}\xi_{2}-\delta_{\xi_{1}}\xi_{2}+\delta_{\xi_{2}}\xi_{1}~,$ (3)
where, $\delta_{\xi_{i}}\xi_{j}$ represents the change induced in $\xi_{j}[g]$ by the variation $\mathcal{L}_{\xi_{i}}g_{\mu\nu}$, that is, $\delta_{\xi_{1}}\xi_{2}^{\mu}=\xi_{2}^{\mu}[\mathcal{L}_{\xi_{1}}g]$ . The (closed) algebra then works out to be
$\xi(T_{1},\Phi_{1}^{A}),\xi(T_{2},\Phi_{2}^{A})]^{\mu}=\xi_{1}^{\nu}\partial_{\nu}\xi_{2}^{\mu}-\xi_{2}^{\nu}\partial_{\nu}\xi_{1}^{\mu}-\xi_{2}^{\mu}[\mathcal{L}_{\xi_{1}}g]+\xi_{1}^{\mu}[\mathcal{L}_{\xi_{2}}g]~,$ (4)
@@
For spherical black holes and constant surface gravity $\kappa$ the algebra has been worked out [18, 19]. One uses the complex coordinates $\zeta,\bar{\zeta}$ on the sphere, expands $T(x^{A})$ and $\Phi^{A}(x^{A})$ in Laurent modes
@@
This is a semi-direct sum of two commuting copies of the Witt algebra generated by $\mathcal{Y}_{m}$ and $\bar{\mathcal{Y}}_{n}$ which represent the superrotations, and an Abelian current algebra generated by $\mathcal{T}_{mn}$’s which represent the supertranslations. This, for example, has been verified for various C-metrics at both their black hole and acceleration horizons as well as for the (cosmological) de Sitter horizon [19, 21]. One can physically interpret the mode $\mathcal{T}_{00}$ as the generator of rigid translations and therefore can be associated with energy. All other generators, i.e., $\mathcal{Y}_{m}$, $\bar{\mathcal{Y}}_{n}$ and $\mathcal{T}_{mn}$, commute with $\mathcal{T}_{00}$, and they can thus be generators of soft hair on the horizon [18].
Charge Algebras: For the charge algebra, one starts with the charge variation, which can be obtained from the covariant formalism [39, 40, 41]
$\cancel{\delta}Q_{\xi}=\frac{1}{16\pi}\int\left(\mathrm{d}^{d-2}x\right)_{\mu\nu}\sqrt{-g}\left[\xi^{[\nu}\nabla^{\mu]}h-\xi^{[\nu}\nabla_{\sigma}h^{\mu]\sigma}+\xi_{\sigma}\nabla^{[\nu}h^{\mu]\sigma}+\frac{1}{2}h\nabla^{[\nu}\xi^{\mu]}-h^{\sigma[\nu}\nabla_{\sigma}\xi^{\mu]}\right]$ (9)
where $\left(\mathrm{d}^{d-2}x\right)_{\mu\nu}\equiv\frac{1}{(d-2)!}\varepsilon_{\mu\nu\alpha_{1}\alpha_{2}\cdots\alpha_{d-2}}\mathrm{d}x^{\alpha_{1}}\wedge\mathrm{d}x^{\alpha_{2}}\wedge\cdots\wedge\mathrm{d}x^{\alpha_{d-2}}$, $h_{\mu\nu}\equiv\delta g_{\mu\nu}$, $\xi$ is the asymptotic Killing vector, and the slash on $\delta$ represents that the variation might not be integrable. Using the boundary conditions (1), and the Killing vectors (2), one obtains the following dynamical fields [20]:
@@
The integrability of $\delta Q$ will depend on the boundary conditions imposed. This is integrable, for example, for constant surface gravity [18, 19]. The Poisson brackets of the charges in this case work out to be a semi-direct sum of two copies of the Witt and an Abelian current algebra, just the same as their Killing vector algebra above in (LABEL:WittAbelian) [17, 18, 19]. This is very close to the BMS (with superrotations) algebra with slightly different structure constants. Note that the expression for variation of charges in [18, 19] (see, for example, equation (2.32) in [18]) looks slightly different from equation (12). However, they are identical expressions in different notations.
Considering the problem in a rotating frame within the Arnowitt–Deser–Misner (ADM) framework of general relativity, one can introduce a “spin parameter” $s$ and impose suitable boundary conditions to ensure that the boundary term — typically required for a well-defined Hamiltonian formulation — is integrable. Under these conditions, a deformation of the BMS group, known as the “spin-$s$ BMS group,” emerges [20] (see Figure 1). For $s=0$, this recovers the algebra obtained by [18, 19], described above in (LABEL:WittAbelian), while for $s=1$, it yields the BMSd algebra. [^6] See [42] for surface charges on a generic null boundary.
@@
A second set of boundary conditions (without the spin parameter) leads to the Heisenberg algebra for the charge generators [20]. The boundary conditions for obtaining the BMS and the Heisenberg algebra are detailed below. For more details, see [20, 44].
Boundary Conditions for BMS: The ADM formalism is characterized by the lapse function $N(x^{\mu})$ and the shift vector $N^{i}(x^{\mu})$. Near a horizon they can be expanded as follows [20]
@@
This factor is crucial in obtaining the spin-$s$ BMS algebra. We will return to this point in section 4 when discussing thermodynamics. The total Hamiltonian is the generator of unit time translations, given by [20]
@@
The Poisson brackets of the charges work out to be [20]
@@
This algebra is the semi-direct sum of diffeomorphisms at the spacelike section of the horizon generated by $\mathcal{J}_{A}$, and an Abelian algebra generated by $\mathcal{P}^{(s)}$. As before, introducing complex coordinates on the sphere, [20] expanded supertranslations and superrotations in Laurent modes. The corresponding generators can be defined as
@@
whose algebra works out to be [20]
$\begin{gathered}\{\mathcal{J}_{k},\mathcal{P}^{(s)}_{(m,n)}\}=\left(\frac{s}{2}(k+1)-m\right)\mathcal{P}^{(s)}_{(k+m,n)}~,\qquad\{\bar{\mathcal{J}}_{k},\mathcal{P}^{(s)}_{(m,n)}\}=\left(\frac{s}{2}(k+1)-n\right)\,\mathcal{P}^{(s)}_{(m,k+n)}\\ \{\mathcal{J}_{n},\mathcal{J}_{m}\}=(n-m)\,\mathcal{J}_{n+m}~,\qquad\{\bar{\mathcal{J}}_{n},\bar{\mathcal{J}}_{m}\}=(n-m)\,\bar{\mathcal{J}}_{n+m}~,\\ \{\mathcal{J}_{n},\,\bar{\mathcal{J}}_{m}\}=\{\mathcal{P}^{(s)}_{(m,n)},\,\mathcal{P}^{(s)}_{(m^{\prime},n^{\prime})}\}=0~.\end{gathered}$ (23)
The third and the fourth brackets — $\{\mathcal{J}_{n},\mathcal{J}_{m}\}$ and $\{\bar{\mathcal{J}}_{n},\bar{\mathcal{J}}_{m}\}$ — denote two copies of the Witt algebra. This is the spin-$s$ BMSd algebra. For $s=0$, this is the algebra for constant surface gravity in [17, 18], which is the Poisson bracket counterpart of (LABEL:WittAbelian). The different BMS groups and their relations are shown in Figure 1.
Figure 1: Evolution of the BMS group: Originally discovered by Bondi, Metzner, van der Burg, and Sachs, the BMS group was later generalized in 2011 by Barnich and Troessaert to include superrotations. Various deformations of the BMS algebra were subsequently studied (see [43] and references therein).
@@
One gets spin-$s$ BMS3 algebra [45, 20]:
@@
Boundary Conditions for Heisenberg-like Algebra: Algebra depends on the choice of slicing used in the solution space [44]. Interchanging the roles of charges $(\mathcal{P},\mathcal{J}_{A})$ and vectors $(T,\Phi^{A})$ in the solution space
@@
the charge variation can also be integrated [20]. Define $\mathcal{J}^{H}_{A}=\mathcal{J}_{A}\mathcal{P}^{-1}$. Then the charges (13) take the form
@@
When $F_{AB}=0$, this is exactly Heisenberg algebra [20].